OpenStudy (anonymous):

Can someone tell me what's the most advanced thing you know in any field of mathematics?

OpenStudy (amistre64):

addition

myininaya (myininaya):

I think some people might find some knowledge more advance then others. i think the word advance is almost an opinion.

hero (hero):

nothing

myininaya (myininaya):

for instance addition is advanced for amistre, but hero finds addition to be not advanced

hero (hero):

Well, it isn't ;P

myininaya (myininaya):

and see amistre would disagree with you

hero (hero):

You still haven't told us your advanced knowledge of math Ms Computational

myininaya (myininaya):

it does depend on what we are adding

myininaya (myininaya):

what if we adding an infinite amount of numbers together

OpenStudy (jamesj):

I know some rather beautiful things in algebraic topology and non-Riemann manifolds. Why?

hero (hero):

In that case the sum is \[ \infty\]

myininaya (myininaya):

sometimes the sum is convergent though

hero (hero):

That's 8 sideways

OpenStudy (anonymous):

I just want to see some cool equations

OpenStudy (anonymous):

What's your favorite equation. Like if you were to show someone 1 equation so that they would be impressed by how cool and advanced it looked, what would it be?

myininaya (myininaya):

\[e^{i \theta}=\cos(\theta)+i \sin(\theta)\]

hero (hero):

I've seen cooler ones

OpenStudy (jamesj):

Hmm. Sometimes big complicated things are not that interesting. But simple looking things can often be deep. For example, this equation is rather thorny. Let p be positive integer greater than 2. Then what are the characteristics of \[x^p + y^p = z^p\]

myininaya (myininaya):

\[curvelength=\int\limits_{a}^{b}\sqrt{1+(f')^2} dx\]